OpenStudy (anonymous):

X squared + 9 = 6x how do i star off

OpenStudy (anonymous):

What are you trying to accomplish?

OpenStudy (anonymous):

direction said to solve that equation by factorying

OpenStudy (anonymous):

factoring

OpenStudy (anonymous):

You will need to have all of your terms on one side of the equation before you begin to factor.

OpenStudy (anonymous):

x^2-6x+9=0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then solve it using Baskara...

OpenStudy (anonymous):

x^2 -6x + 9 = 0 X^2 -3x -3x + 9 = 0 x(x-3) -3(x-3) = 0 (x-3)(x-3) = 0 (x-3)^2 = 0

OpenStudy (anonymous):

4x^2-13x+12=0

OpenStudy (anonymous):

do you mean 4x^2-14x+12=0, if not the answer is non-reals